# On the Low-Temperature MCMC threshold:

### v2

@@ 8.2. Proof of Positive Results for Sparse Linear Regression
Here we prove Theorem 4.2 under Assumption 4.1. First we show that Assumption 4.1 implies Assumption 2.2.
8.2.1. Verifying Assumption 2.2
In order to verify Assumption 2.2 we present a Binary - Local Search Algorithm that is a binary version of the Local Search Algorithm in [GZ22a]. The proof of its convergence is essentially the same as for the Local Search Algorithm, except we only need to consider support recovery. We use $e_{1},e_{2},\dots$ to refer to the standard basis vectors in $\mathbb{R}^{p}$. We write $X_{i}$ to refer to the $i^{th}$ column of $X$.
@@
B-LSA looks at all the possible coordinate “flips” we could make in one step, and chooses the one that minimises the squared error. It repeats until no local improvement is possible. In the following lemma, we verify the second part of Assumption 2.2 under Assumption 4.1, with $\delta=\frac{n}{4}$ by showing that B-LSA always reduces the squared error by a factor of at least order $n$ at each step.
@@
Assume that Assumption 4.1 holds. If $v\neq v^{*}$, then for any $v^{\prime}$ obtained from $v$ in one iteration of B-LSA, with high probability it holds that
@@
The proof of Lemma 8.4 is deferred until Section 12. In the next lemma, we verify the first part of Assumption 2.2 with $D=O(k)$.
@@
Assume that Assumption 4.1 holds. With high probability as $k\rightarrow\infty$, B-LSA with uniform random initialisation $v_{0}$ terminates in at most $O(k)$ iterations and outputs $v^{*}$.
@@
We will now use the standard $\chi^{2}$ bound (for instance, page 1325 of [LM00]) that
@@
Note that since $\sigma^{2}\leq ck$ by Assumption 4.1, we have that $20\left(2(k-\ell)+\sigma^{2}\right)\leq(40+20c)k$, and therefore
$\mathbb{P}\left(\frac{4}{n}\lVert Y-Xv_{0}\rVert_{2}^{2}\geq(40+20c)k\right)=o(1)$
@@
We assume Assumption 4.1 and $n\geq Ck\log(p/k)$ for some $C>0.$ Then with high probability as $k\rightarrow\infty$, it holds that
@@
and thus we get that with high probability (and recalling that $\sigma^{2}\leq ck$)
@@
Note that since $n=\Omega(k\log(p/k)$ and $\sigma^{2}\leq ck$, all terms are $O(kn)$, and the result follows.
$\max_{v}\lVert Y-Xv\rVert_{2}^{2}=O(kn)$
Therefore Assumption 2.2 is satisfied with $D=O(k)$ and $\delta=\frac{n}{4}$, and $\mathcal{R}_{H}=O(kn).$ We now have everything we need to prove Theorem 4.2.
@@
We note that in this case that $\Delta$, the maximum degree of the graph, is given by $\Delta=\max_{v\in\mathcal{X}_{p,k}}|\Gamma_{G}(v)|=k(p-k)$, as each $v$ has $k$ 1’s and $(p-k)$ 0’s. By Lemma 8.4, Lemma 8.5, and Lemma 8.6, we have that Assumption 2.2 holds with $D=O(k)$, $\delta=\frac{n}{4}$, and $\mathcal{R}_{H}=O(kn)$. We also see that the cardinality of the state space $\mathcal{X}$ is given by $\binom{p}{k}$. We therefore apply the first two parts of Theorem 2.3 to see that if
@@
iterations, where we use $\binom{p}{k}\leq\left(\frac{pe}{k}\right)^{k}$. For the second part of Theorem 4.2, we apply the final part of Theorem 2.3. This part requires
@@
Note that since $n\geq Ck\log(\frac{p}{k})$, $\frac{8}{n}k\log\left(\frac{pe}{k}\right)$ is at most $8/C$ plus an $o(1)$ term. For the second term: $\mathcal{R}_{H}=O(kn)$, so $\frac{8}{n}\log\left(\frac{160\mathcal{R}_{H}}{n}\right)=O\left(\frac{\log k}{n}\right)$, which is dominated by the first term since $n=\omega(k)$. Therefore the condition on $\beta$ is satisfied whenever $\beta$ is larger than $8/C$. It then follows by Theorem 2.3 that the Metropolis chain finds $v^{*}$ within
@@
iterations. ∎
@@ 9. Proofs of Negative Results for the Gaussian Additive Model
Here we show the tightness of our positive results for the Gaussian Additive Model by proving Theorem 3.6, a superpolynomial lower bound on the mixing time for Sparse PCA. The general approach we follow for both this lower bound, and the lower bound in sparse regression that follows, is a careful implementation of the “Overlap Gap Property for inference” framework from [GZ22a], which is inspired by statistical physics. This Overlap Gap Property framework has been shown in previous works (e.g., in [GJS21, GZ19, AWZ23]) to directly imply mixing time lower bounds.
In this work though, to keep the presentation simple and focused on Markov chains, we avoid defining exactly what the Overlap Gap Property is, but follow only the proof steps required to derive directly the desired mixing lower bound. We direct the reader to [AWZ23] for the more details on the underlying connections.
@@ 9.1. Auxiliary Lemmas
We start with two folklore inequalities, see e.g., [Sav62].
@@
Let $q=\left\lfloor k^{2}/p\right\rfloor$ denote the “typical” overlap of a random $k$-sparse binary vector $x$ with the hidden signal $v$. Note that $q=0$ if $k<\sqrt{p}$, and $q=o(k)$ since $k=o(p)$. We will show that
@@ 9.2. Proof Strategy
Let us consider the following key random functions (analogues of the key quantities defined in [GZ22a]) which control the maximum log-likelihood value that can be achieved by any binary $k$-sparse $x$ of overlap $l=0,1,\ldots,k$ with $v^{*}$:
@@
In [GZ22a] the authors use the non-monotonicity of $\Gamma_{\ell}$ to obtain the MCMC lower bound. To achieve a similar result we need to understand the concentration properties of $\Gamma_{\ell}$.
@@
$\log\left(\binom{k}{\ell}\binom{p-k}{k-\ell}\right)\geq\log\binom{p-k}{k-\ell}\geq(1-o(1))k\log(p/k).$
@@
and in particular $\Gamma_{\ell}$ must be strictly decreasing for some $\ell\in[0,\ell^{*}].$ We remark this is the key property which leads to the presence of the Ovelap Gap Property as in [GZ22a].
@@
Theorem 3.6 then follows from [MWW09, Claim 2.1] (see also [AWZ20, Proposition 2.2]).
@@ 9.4. The case $t=2$
$\frac{1}{\ell}s^{\intercal}Ws=\frac{\rho}{k}x^{\intercal}Wx+\sqrt{1-\rho^{2}}Z_{s},$
@@
$\left|s^{\intercal}Ws-\frac{\ell^{2}}{k^{2}}x^{\intercal}Wx\right|\leq\frac{2\ell}{k}\sqrt{(k^{2}-\ell^{2})\log\binom{k}{\ell}}\leq 2\sqrt{(k^{2}-\ell^{2})\log\binom{k}{\ell}},$
@@ 10. Proof of Negative Results for Sparse Regression
In this section, we prove Theorem 4.4, the lower bound on the mixing time for sparse regression. This shows the tightness of our positive results for sparse regression. We do this by a similar but significantly improved approach to prove the Overlap Gap Property result in [GZ22a] which then implies the mixing time lower bound. As in the previous section, we omit the Overlap Gap Property specifics as much as possible, and focus on the desired mixing time lower bound.
@@ 10.1. Improving Previous Results
Towards proving our negative result, we first need to generalize some key results from [GZ22a]. We consider the following two constrained optimization problems:
@@
We now present strengthened versions of two theorems from [GZ22a]. Both of these have been strengthened to assume $k=o(p^{1/3-\eta})$ for some $\eta>0$ and no assumption on $n$, instead of the more restrictive $k\log k=O(n)$ in [GZ22a], which requires $k=p^{o(1)}$ in the relevant regime $n=o(k\log(p/k)).$
@@
[Improved version of Theorem 2.1 of [GZ22a]] Suppose for some $\eta>0$, $k=o(p^{1/3-\eta})$. Then:
@@
The proof of Theorem 10.1(a) follows identically to the proof in [GZ22b], but with $k\log(p/k)$ in place of $k\log p$ inside the exponent. The proof of Theorem 10.1(b) uses a second moment method (which is the bulk of the proof), which we defer to Section 12.
@@ 10.2. Proof of the main negative result
It holds that
@@
Let $v\in A_{1}$. Then by Lemma 10.2 and Lemma 10.3,
@@
Using Theorem 10.5, Theorem 4.4 follows from [MWW09, Claim 2.1] (see also [AWZ20, Proposition 2.2]).
@@ 11. Computational hardness of sparse tensor PCA
$\lambda_{ALG}(t,k)=\tilde{\Theta}(\min\{k^{t/2},p^{(t-1)/2}/k^{t/2-1}\}).$
The argument is essentially borrowed from a more general argument in the proof of [LZ22, Theorem 16]. Yet, as we are only interested in a specific implication of the proof of [LZ22], not formally covered by any of the theorems in [LZ22], we describe it here for reasons of completeness and the reader’s convenience.
@@ 11.1. The hypergraph planted dense subgraph problem
In HPDS, given $n,k=k_{p}\in\mathbb{N},k\leq p$, some $\rho=\rho_{p}\in(0,1)$ and some fixed $t\geq 2,$ one observes an instance of a random $t$-hypergraph on $p$ vertices, which we denote by $G$ and is constructed as follows. Firstly, one chooses $k$ vertices out of $n$ uniformly at random, resulting in a $k$-subset $\mathcal{PC}$. Then, given $\mathcal{PC},$ $G$ is a random $t$-hypergraph which contains all $t$-hyperedges between vertices of $\mathcal{PC}$ independently with probability $1/2+\rho$, and all other $t$-hyperedges independently and with probability exactly $1/2$ The goal of the statistician is to recover $\mathcal{PC}$ from observing $G$ with high probability as $p\rightarrow+\infty$.
Of course the larger the “gap parameter” $\rho\in(0,1)$ is, the easier the recovery of $\mathcal{PC}$ from $G$ is. In terms of computational limits for HPDS, a direct application of [LZ22, Theorem 20] implies that for some
$\rho_{\mathrm{ALG}}=\tilde{\Theta}(\min\{1,p^{(t-1)/2}/k^{t-1}\}),$
if $\rho=o(\rho_{\mathrm{ALG}})$ then all $(\log p)^{2}$-degree polynomials fail to recover $\mathcal{PC}$ from $G$. This result provides strong evidence of computational hardness in that regime of HPDS, based on the low-degree framework for estimation problems [SW22, LZ22].
@@ 11.2. Connecting back to sparse tensor PCA
$\lambda_{ALG}(t,k)=\tilde{\Theta}(k^{t/2}\log(1+2\rho_{\mathrm{ALG}}))=\tilde{\Theta}(k^{t/2}\rho_{\mathrm{ALG}})=\tilde{\Theta}(\min\{k^{t/2},p^{(t-1)/2}/k^{t/2-1}\}),$
@@
The simple reduction works by employing the rejection kernel technique [BBH18]. Indeed by using the technique as described in [LZ22, Lemma 15] for probabilities 1/$21/2+\rho,1/$ 2 and some $\xi=\tilde{\Theta}(\log(1+2\rho)/\sqrt{t})$, combined with the tensorization of the total variation distance [LZ22, Lemma 13], we directly get that the total variation between an instance $G$ from HPDS with parameter $\rho$ and an instance $Y$ from sparse tensor PCA for $\lambda=k^{t/2}\xi$ is $o(1)$, which completes the reduction.
12. Deferred proofs
12.1. Proof of Lemma 8.4
Here we prove Lemma 8.4, that the Binary - Local Search Algorithm decreases the squared error by at least $\frac{1}{4}n$ at every step when $n\geq Ck\log(p/k)$ and $\sigma^{2}\leq c\min\left\{\frac{\log p}{\log\log p},k\right\}$. The proof is essentially identical to the proof of Theorem 2.12 in [GZ22a], but applied to the binary case. We will need a couple of results from [GZ22b].
Definition 12.1.

### v3

@@ 8.2. Proof of Positive Results for Sparse Linear Regression
Here we prove Theorem 4.2 under 4.1. First we show that 4.1 implies 2.2.
8.2.1. Verifying 2.2
In order to verify 2.2 we present a Binary - Local Search Algorithm that is a binary version of the Local Search Algorithm in [26]. The proof of its convergence is essentially the same as for the Local Search Algorithm, except we only need to consider support recovery. We use $e_{1},e_{2},\dots$ to refer to the standard basis vectors in $\mathbb{R}^{p}$. We write $X_{i}$ to refer to the $i^{th}$ column of $X$.
@@
B-LSA looks at all the possible coordinate “flips” we could make in one step, and chooses the one that minimises the squared error. It repeats until no local improvement is possible. In the following lemma, we verify the second part of 2.2 under 4.1, with $\delta=\frac{n}{4}$ by showing that B-LSA always reduces the squared error by a factor of at least order $n$ at each step.
@@
Assume that 4.1 holds. If $v\neq v^{*}$, then for any $v^{\prime}$ obtained from $v$ in one iteration of B-LSA, with high probability it holds that
@@
The proof of Lemma 8.4 is deferred to Section A.1. In the next lemma, we verify the first part of 2.2 with $D=O(k)$.
@@
Assume that 4.1 holds. With high probability as $k\rightarrow\infty$, B-LSA with uniform random initialisation $v_{0}$ terminates in at most $O(k)$ iterations and outputs $v^{*}$.
@@
We will now use the standard $\chi^{2}$ bound (for instance, page 1325 of [39]) that
@@
Note that since $\sigma^{2}=o(k)$ by 4.1, we have that $20\left(2(k-\ell)+\sigma^{2}\right)\leq 50k$ for large enough $k$, and therefore
$\mathbb{P}\left(\frac{4}{n}\lVert Y-Xv_{0}\rVert_{2}^{2}\geq 50k\right)=o(1)$
@@
We assume 4.1 and $n\geq Ck\log(p/k)$ for some $C>0$. Then with high probability as $k\rightarrow\infty$, it holds that
@@
and thus we get that with high probability (and recalling that $\sigma^{2}=o(k)$),
@@
Note that since $n=\Omega(k\log(p/k)$ and $\sigma^{2}=o(k)$, all terms are $O(kn)$, and the result follows. ∎
Therefore 2.2 is satisfied with $D=O(k)$ and $\delta=\frac{n}{4}$, and $\mathcal{R}_{H}=O(kn)$. We now have everything we need to prove Theorem 4.2.
@@
We note that in this case that $\Delta$, the maximum degree of the graph, is given by $\Delta=\max_{v\in\mathcal{X}_{p,k}}|\Gamma_{G}(v)|=k(p-k)$, as each $v$ has $k$ 1’s and $(p-k)$ 0’s. By Lemma 8.4, Lemma 8.5, and Lemma 8.6, we have that 2.2 holds with $D=O(k)$, $\delta=\frac{n}{4}$, and $\mathcal{R}_{H}=O(kn)$. We also see that the cardinality of the state space $\mathcal{X}$ is given by $\binom{p}{k}$. We therefore apply the first two parts of Theorem 2.3 to see that if
@@
iterations, where we use $\binom{p}{k}\leq\left(\frac{pe}{k}\right)^{k}$. This proves the moderate $\beta$ statement. For the large $\beta$ statement, we apply the final part of Theorem 2.3. This part requires
@@
Note that since $n\geq Ck\log(\frac{p}{k})$, $\frac{8}{n}k\log\left(\frac{pe}{k}\right)$ is at most $8/C$ plus an $o(1)$ term. For the second term: $\mathcal{R}_{H}=O(kn)$, so $\frac{8}{n}\log\left(\frac{160\mathcal{R}_{H}}{n}\right)=O\left(\frac{\log k}{n}\right)$, which is dominated by the first term since $n=\omega(k)$. Therefore the condition on $\beta$ is satisfied whenever $\beta$ is larger than $8/C$. It then follows by Theorem 2.3 that the Metropolis chain after
@@
iterations is at $v^{*}$ with probability at least $2/3-o(1)$. Finally, we note that by Lemma 8.4 that for any neighbor $v$ of $v^{*}$, it holds with high probability that $H(v^{*})-H(v)\geq n/4$. Therefore, with high probability, the Randomized Greedy algorithm (i.e. $\beta=\infty$) never leaves $v^{*}$ once it hits it. ∎
@@ 9. Proofs of Negative Results for the Gaussian Additive Model
Here we show the tightness of our positive results for the Gaussian Additive Model by proving Theorem 3.8, a superpolynomial lower bound on the mixing time for Sparse PCA. The general approach we follow for both this lower bound, and the lower bound in sparse regression that follows, is a careful implementation of the “Overlap Gap Property for inference” framework from [26], which is inspired by statistical physics. This Overlap Gap Property framework has been shown in previous works (e.g., in [24, 28, 2]) to directly imply mixing time lower bounds.
In this work though, to keep the presentation simple and focused on Markov chains, we avoid defining exactly what the Overlap Gap Property is, but follow only the proof steps required to derive directly the desired mixing lower bound. We direct the reader to [2] for the more details on the underlying connections.
@@ 9.1. Auxiliary Lemmas
We start with two folklore inequalities, see e.g., [58].
@@
Let $q=\left\lfloor k^{2}/p\right\rfloor$ denote the “typical” overlap of a random $k$-sparse binary vector $x$ with the hidden signal $v^{*}$. Note that $q=0$ if $k<\sqrt{p}$, and $q=o(k)$ since $k=o(p)$. We will show that
@@ 9.2. Proof Strategy
Let us consider the following key random functions (analogues of the key quantities defined in [26]) which control the maximum log-likelihood value that can be achieved by any binary $k$-sparse $x$ of overlap $l=0,1,\ldots,k$ with $v^{*}$:
@@
In [26] the authors use the non-monotonicity of $\Gamma_{\ell}$ to obtain the MCMC lower bound. To achieve a similar result we need to understand the concentration properties of $\Gamma_{\ell}$.
@@
$\log\left(\binom{k}{\ell}\binom{p-k}{k-\ell}\right)\geq\log\binom{p-k}{k-\ell}\geq(1-o(1))k\log(p/k)$
@@
and in particular $\Gamma_{\ell}$ must be strictly decreasing for some $\ell\in[0,\ell^{*}]$. We remark this is the key property which leads to the presence of the Ovelap Gap Property as in [26].
@@
Theorem 3.8 then follows from [52, Claim 2.1] (see also [1, Proposition 2.2]).
@@ 9.4. The case $t=2$
$\frac{1}{\ell}s^{\intercal}Ws=\frac{\rho}{k}x^{\intercal}Wx+\sqrt{1-\rho^{2}}Z_{s}$
@@
$\left|s^{\intercal}Ws-\frac{\ell^{2}}{k^{2}}x^{\intercal}Wx\right|\leq\frac{2\ell}{k}\sqrt{(k^{2}-\ell^{2})\log\binom{k}{\ell}}\leq 2\sqrt{(k^{2}-\ell^{2})\log\binom{k}{\ell}}$
@@ 10. Proof of Negative Results for Sparse Regression
In this section, we prove Theorem 4.5, the lower bound on the mixing time for sparse regression. This shows the tightness of our positive results for sparse regression. We do this by a similar but significantly improved approach to prove the Overlap Gap Property result in [26] which then implies the mixing time lower bound. As in the previous section, we omit the Overlap Gap Property specifics as much as possible, and focus on the desired mixing time lower bound.
@@ 10.1. Improving Previous Results
Towards proving our negative result, we first need to generalize some key results from [26]. We consider the following two constrained optimization problems:
@@
We now present strengthened versions of two theorems from [26]. Both of these have been strengthened to assume $k=o(p^{1/3-\eta})$ for some $\eta>0$ and no assumption on $n$, instead of the more restrictive $k\log k=O(n)$ in [26], which requires $k=p^{o(1)}$ in the relevant regime $n=o(k\log(p/k))$.
@@
[Improved version of Theorem 2.1 of [26]] Suppose for some $\eta>0$, $k=o(p^{1/3-\eta})$. Then:
@@
The proof of Theorem 10.1(a) follows identically to the proof in [27], but with $k\log(p/k)$ in place of $k\log p$ inside the exponent. The proof of Theorem 10.1(b) uses a second moment method (which is the bulk of the proof), which we defer to Section A.2.
@@ 10.2. Proof of the main negative result
With high probability, it holds that
@@
Let $v\in A_{1}$. Then by Lemma 10.2 and Lemma 10.3, with high probability we have
@@
Using Theorem 10.5, Theorem 4.5 follows from [52, Claim 2.1] (see also [1, Proposition 2.2]).
@@ 11. Computational hardness of sparse tensor PCA
$\lambda_{ALG}(t,k)=\tilde{\Theta}(\min\{k^{t/2},p^{(t-1)/2}/k^{t/2-1}\})$
The reduction is based on a similar reduction described in [10, Theorem 10] for where $v^{*}\in\{-1,1\}^{n}$ without sparsity, which we appropriately extend to the case of a $k$-sparse binary $v^{*}$. Our modification of the argument given here is essentially borrowed from a more general argument described in the proof of [44, Theorem 16]. Yet, as we are only interested in a specific implication of the proof in [44], not formally covered by any of the theorems in [44] or [10], we describe it here for reasons of completeness and the reader’s convenience.
@@ 11.1. The hypergraph planted dense subgraph problem
In HPDS, given $n,k=k_{p}\in\mathbb{N},k\leq p$, some $\rho=\rho_{p}\in(0,1)$ and some fixed $t\geq 2$, one observes an instance of a random $t$-hypergraph on $p$ vertices, which we denote by $G$ and is constructed as follows. Firstly, one chooses $k$ vertices out of $n$ uniformly at random, resulting in a $k$-subset $\mathcal{PC}$. Then, given $\mathcal{PC}$, $G$ is a random $t$-hypergraph which contains all $t$-hyperedges between vertices of $\mathcal{PC}$ independently with probability $1/2+\rho$, and all other $t$-hyperedges independently and with probability exactly $1/2$. The goal of the statistician is to recover $\mathcal{PC}$ from observing $G$ with high probability as $p\rightarrow+\infty$.
Of course the larger the “gap parameter” $\rho\in(0,1)$ is, the easier the recovery of $\mathcal{PC}$ from $G$ is. In terms of computational limits for HPDS, a direct application of [44, Theorem 20] implies that for some
$\rho_{\mathrm{ALG}}=\tilde{\Theta}(\min\{1,p^{(t-1)/2}/k^{t-1}\})$
if $\rho=o(\rho_{\mathrm{ALG}})$ then all $(\log p)^{2}$-degree polynomials fail to recover $\mathcal{PC}$ from $G$. This result provides strong evidence of computational hardness in that regime of HPDS, based on the low-degree framework for estimation problems [59, 44].
@@ 11.2. Connecting back to sparse tensor PCA
$\lambda_{ALG}(t,k)=\tilde{\Theta}(k^{t/2}\log(1+2\rho_{\mathrm{ALG}}))=\tilde{\Theta}(k^{t/2}\rho_{\mathrm{ALG}})=\tilde{\Theta}(\min\{k^{t/2},p^{(t-1)/2}/k^{t/2-1}\})$
@@
The simple reduction works by employing the rejection kernel technique [11]. Indeed by using the technique as described in [44, Lemma 15] for probabilities 1/$21/2+\rho,1/$ 2 and some $\xi=\tilde{\Theta}(\log(1+2\rho)/\sqrt{t})$, combined with the tensorization of the total variation distance [44, Lemma 13], we directly get that the total variation between an instance $G$ from HPDS with parameter $\rho$ and an instance $Y$ from sparse tensor PCA for $\lambda=k^{t/2}\xi$ is $o(1)$, which completes the reduction.
12. Acknowledgments
Appendix A Deferred Proofs
A.1. Proof of Lemma 8.4
Here we prove Lemma 8.4, that the Binary - Local Search Algorithm decreases the squared error by at least $\frac{1}{4}n$ at every step when $n\geq Ck\log(p/k)$ and $\sigma^{2}\leq k/\log p$. The proof is based on an appropriately strengthened version of the proof of Theorem 2.12 in [26], when applied to the case of binary coefficients. We will need a couple of results from [27].
Definition A.1.